OpenStudy (anonymous):

Fluffy, Spot, and Shampy have a combined age in dog years of 91. Spot is 14 years younger than Fluffy. Shampy is 6 years older than Fluffy. What is Fluffy's age, f, in dog years? (Change the following question to an algebraic equations.)

OpenStudy (anonymous):

X+X-14+X+6=91

OpenStudy (anonymous):

but when you combine like terms you get 3x-12=91

OpenStudy (anonymous):

the question iis to make an algebraic equesion to the quetion and that is how you do it for this kind of question

OpenStudy (anonymous):

-14+6=-8

OpenStudy (anonymous):

and then 3x=103 and then x=103/3 which is like 34.33333333333 but yeah i guess youre right for all purposes of an equation

OpenStudy (anonymous):

3x-8=91 3x=99 x=33

OpenStudy (anonymous):

sorry -14+6=-8 not -12

OpenStudy (anonymous):

ohhhhhh thanks hahaha i hate when i mess up simple math like that